OpenStudy (anonymous):

square root of x^3 divided by 81x^5y^6

OpenStudy (anonymous):

\[\sqrt{\frac{\cancel x*x^2}{81*x^2*x^2*\cancel x*y^2*y^2*y^2}}=\frac{x}{9x^2y^3}\]